# Physics-informed Mesh-independent Deep Compositional Operator Network

### v1

@@ Abstract
Solving parametric Partial Differential Equations (PDEs) for a broad range of parameters is a critical challenge in scientific computing. To this end, neural operators, which learn mappings from parameters to solutions, have been successfully used. However, the training of neural operators typically demands large training datasets, the acquisition of which can be prohibitively expensive. To address this challenge, physics-informed training can offer a cost-effective strategy. However, current physics-informed neural operators face limitations, either in handling irregular domain shapes or in generalization to various discretizations of PDE parameters with variable mesh sizes. In this research, we introduce a novel physics-informed model architecture which can generalize to parameter discretizations of variable size and irregular domain shapes. Particularly, inspired by deep operator neural networks, our model involves a discretization-independent learning of parameter embedding repeatedly, and this parameter embedding is integrated with the response embeddings through multiple compositional layers, for more expressivity. Numerical results demonstrate the accuracy and efficiency of the proposed method.
Keywords Physics-informed neural networks, Neural operators, Mesh generalization
@@ 1 Introduction
Recently, various neural operator architectures have been proposed for operator learning in a data-driven (supervised learning) approach Lu et al. (2021); Li et al. (2020a); Kovachki et al. (2021); Li et al. (2020b). As the pioneering work in this area, Deep Operator Networks (DeepONet) Lu et al. (2021) were developed, based on an inspiration from the universal approximation theorem for operators. DeepONet approximates nonlinear operators by effectively learning a collection of basis functions and coefficients, in the form of neural networks. However, the function input to DeepONet, i.e. the PDE parameters, are structured in the form of a finite-dimensional vector with a fixed dimension. This makes DeepONet to be a "mesh-specific" method Kovachki et al. (2021). For instance, consider employing sensors on $m$ distinct locations to capture a observation of the PDE parameter, which gives us a $m$-dimensional representation of the PDE parameters. Then a DeepONet is trained as a mapping from an $m$-dimensional vector input, and cannot be used for cases where more or fewer sensors are used, or to discretizations with a different size. This restriction narrows the utility of neural operators for practical engineering challenges, where various mesh sizes may be used.
@@
To overcome the limitations of DeepONet, numerous "mesh-independent" neural operators have been introduced. The Graph Neural Operator (GNO) Li et al. (2020b) adopts graph neural networks (GNNs) architecture for operator learning, i.e. by considering graph structures for model’s input and output. The Fourier Neural Operator (FNO) Li et al. (2020a) leverages the Fourier transform to learn mappings in the spectral domain, capturing global dependencies and showing superior performance in rectangular domain shapes. Inspired by FNO, Wavelet Neural Operator (WNO) Tripura and Chakraborty (2022) uses the Wavelet Transformation, instead of the Fourier transformation, to better model signals with discontinuity and spikes, and also handle non-square domains. Also, the Low-Rank Neural Operator (LRNO) Kovachki et al. (2021) boosts the efficiency and scalability of neural operator models, using a low-rank representation of the kernels in neural operators.
@@
While these "mesh-independent" methods have shown promise in various applications, they all are data-intensive Wang et al. (2021). Training neural operators in a data-driven (supervised) way necessitates collection of a large number of parameter-solution pairs as the training set, which is obtained using repeated costly high-fidelity (FEM) simulations. This is prohibitively expensive in complex industrial applications. Under this circumstance, physics-informed neural operators emerge as an effective solution for constructing neural operators. These methods draw inspiration from physics-informed neural networks (PINNs) Raissi et al. (2019), and integrate the governing PDEs directly into the training process. This allows for a data-free training without need for FEM-based training data. Among the first physics-based methods, parametric physics-informed neural networks Zhu et al. (2023); Beltrán-Pulido et al. (2022); Naderibeni et al. (2024) considered various elements of a PDE problem, such as PDE parameters Beltrán-Pulido et al. (2022); Hansen et al. (2023), boundary conditions Zhu et al. (2023), and even the geometry of the domain Naderibeni et al. (2024), as additional feature vectors. These feature vectors provide additional inputs to the model, allowing it to learn the dependence of the PDE solution on PDE parameters. However, this approach is limited to cases where an analytical representation of varying conditions, such as boundary conditions and geometries, are available, and cannot be effective where these conditions are available in discrete form with many nodes.
@@
Another class of data-free methods are neural operators that are trained using a physics-informed loss function Goswami et al. (2023); Ren et al. (2022); Moseley et al. (2023); Baddoo et al. (2023). Among them, physics-informed DeepONet (PI-DeepONet) Wang et al. (2021) extends the original DeepONet framework by embedding physical laws directly into the loss function which can be effectively calculated in any irregular domain shape. The incorporation of physical laws also enhances the learning process but, even with this advancement, physics-informed DeepONet cannot generalize to parameter discretizations with different size. On the other hand, the physics-informed Fourier Neural Operator (PI-FNO) Li et al. (2021b) builds upon the original FNO framework, and offers generalization across different parameter dicretizations while also incorporating physics-informed training methods. However, the the reliance of PI-FNO on Fast Fourier Transform (FFT) restricts its application to rectangular domains with uniform meshes as FFT works on uniform sampling of the signal in the spatial domain Li et al. (2023). Another notable development is the physics-informed Wavelet Neural Operator (PI-WNO) Navaneeth et al. (2024), which implements physics-informed training for WNO Tripura and Chakraborty (2022) using stochastic projection. Although PI-WNO can handle various PDE parameter discretizations, the hidden kernels are learned based on the architecture of convolution neural networks O’Shea and Nash (2015), which prevents the model from learning in complex domain geometries.
@@
In our study, we tackle these related due to data requirement, discretization-dependence and irregular geometries by introducing an innovative neural operator architecture. Our Physics-informed Deep Compositional Operator Network (PI-DCON) model is capable of generalizing across different PDE parameter discretizations, including those in irregular domain shapes. In this study, the irregular domain shapes are defined as those that does not belong to regular geometric forms such as rectangles, circles, or polygons. The differences between our proposed model versus other existing works are summarized in Table LABEL:table.archi_compare. To the best of our knowledge, our paper is the first attempt to develop a mesh-independent neural operator that can handle complex geometries without need for any training data.
@@ 3 Methodology
Without loss of generality, let us consider a PDE operator mapping from boundary conditions to the PDE solution, denoted by $\mathcal{M}:g(\bm{x}^{\prime})\rightarrow u(\bm{x})$. In DCON, similarly to DeepONet, the boundary conditions is encoded using a finite number of sampled boundary conditions $[\bm{x}^{\prime}_{1},\bm{x}^{\prime}_{2},...,\bm{x}^{\prime}_{m}]$, and the corresponding function values $[g(\bm{x}^{\prime}_{1}),g(\bm{x}^{\prime}_{2}),...,g(\bm{x}^{\prime}_{m})]$. The architecture of DCON consists of a "branch network" and a "solution network". The branch network maps the given boundary conditions on each of the available $m$ boundary locations to a higher dimensional space, $\mathcal{S}_{\text{param}}$, using a multi-layer perceptron. Then, a Max-pooling layer is applied to this high-dimensional representation to capture, in each dimension of $\mathcal{S}_{\text{param}}$, the most important feature across the $m$ locations on the boundary. The final output of the branch network is a global parameter embedding $\bm{b}$ in the high-dimensional space $\mathcal{S}_{\text{param}}$, whose dimension is independent of the number of boundary points, $m$. This setup relaxes the requirement for all training and test data to include the same number of boundary locations, $m$.
@@
where $\sigma$ represents the nonlinear activation function, $W_{\bm{t}}\in R^{q\times d},B_{\bm{t}}\in R^{q}$ are trainable parameters of the mapping $\bm{t}$, and $W^{j}_{L}\in R^{q\times q},B^{j}_{L}\in R^{q}$ are trainable parameters of the $j$-th mapping $N$. These parameters of the model are optimized by minimizing the physics-informed training loss as shown in equation 2. More details about the proposed model architecture are shown in Figure 2. It should be noted that the proposed DCON architecture can also be used in a data-driven approach, where its parameters are estimated in a supervised learning process, as will be discussed in Section 4.3.
@@ 4 Numerical results
In this section, we numerically evaluate the accuracy of the proposed model for solving parametric differential equations. We compare our models with data-driven DeepONet Lu et al. (2021) and physics-informed DeepONet Wang et al. (2021). We use these default settings unless mentioned otherwise; We use the hyperbolic tangent function (Tanh) Lau and Lim (2018) as our activation function to ensure the smoothness in high-order derivatives. For the architecture of DeepONet, we use three hidden layers of width 512. For our model architecture, we use three operator layers of width 512. Adam is the default optimizer with the following default hyper-parameters: $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999 Kingma and Ba (2014). Each model is trained on 70% of the sampled PDE parameters and validated on 10% of the sampled PDE parameters. Well-trained models predict the PDE solution for the remaining 20% of the sampled PDE parameters. Two hyper-parameters used in model training: (1) the learning rate, with the possible values of learning rate are 0.001, 0.0005, 0.0002, and 0.0001; and (2) the coordinate sampling size ratio in each epoch. This ratio determines the number of collocation points in each epoch as a function of the number of sampled PDE parameters. The possible values for this ratio are 0.3, 0.2, 0.1, and 0.05. We implement grid search Bergstra and Bengio (2012) to calculate the best hyper-parameters, for which report the corresponding model performances are reported. Model training is performed on an NVIDIA P100 GPU using a batch size of 20. We collect the parameters of PDEs by Monte Carlo Simulation Harrison (2010) of stochastic processes Parzen (1999) and derive the solution of PDEs based on the Finite Element Method using Matlab Matlab (2012).
@@ 4.1.1 A Darcy flow problem
We employ this Gaussian process model of Eq. 13 with $l=1$ to generate $N=1000$ different imposed pressure functions on $\partial\Omega_{2}$. Then given these sampled functions, $\{g_{i}\}_{i=1}^{N}$ we solve the PDE using Finite Element Method (FEM) Li and Chen (2019). In order to demonstrate the resol-independence feature of our model, for different sampled boundary functions, we consider different mesh sizes in the domain. Specifically, for the $i$-th realization, with the boundary condition function $g_{i}$, let $M_{1,i}$ and $M_{2,i}$ be the number of sampled locations on $\partial\Omega_{1}$ and $\partial\Omega_{2}$, respectively, and $\bm{X}_{\text{BC}_{1},i}$ and $\bm{X}_{\text{BC}_{2},i}$ are the corresponding sets of coordinates at those sampled locations, i.e.
@@ 4.1.2 A 2D plate problem
where $u$ and $v$ are the plate displacements in $x$ and $y$ directions, respectively, $E$ is the Young’s Modules, and $\mu$ is the Poisson’s Ratio.
The plate is 20 mm × 20 mm, and the hole has a diameter of 5 mm. We impose different boundary conditions on the different edges of the plate. The edge surrounding the central hole is subjected to a fixed boundary condition and there are imposed displacements on the left and right sides. The top and bottom sides are assigned free boundary conditions. The complete boundary conditions are shown in Figure  4, and are given by
@@ 4.2 Main results
For each problem, we conduct two experiments. First, in order to compare our method with PI-DeepONet, which can only handle data with the same mesh size, we consider a fixed mesh size, taken to be small, for all the training and test cases. In the second experiment, we relax this constraint and generate training and test data each with a variable mesh size. The comparison between our proposed methods and the PI-DeepONet is summarized in Table  LABEL:fig.PI_com. We use the relative $L_{2}$ error to evaluate the accuracy of models. We report the average error calculated across the entire test dataset, along with the standard deviation of the relative errors, to demonstrate the model stability. Our proposed model showed satisfactory mean performance in both the Darcy flow problem and the 2D plate problem. Compared with physics-informed DeepONet, our model achieved 64.8% accuracy improvement for the Darcy flow problem and 68.1% accuracy improvement for the 2D plate problem. Based on the standard deviation of relative errors, our model also shows higher stability.
In the second experiment, when variable mesh sizes are used, compared to the fixed mesh size experiment, we note that PI-DCON shows reduction in accuracy for both the Darcy flow and the 2D plate problems. These results are reasonable, especially considering the 2D plate problem is inherently more complex. Despite these challenges, our model is able to maintain acceptable levels of accuracy, with mean relative errors below 3.5% for both problems.
Table 2: Accuracy comparison between PI-DCON and PI-DeepONet.
Mean of relative error Standard deviation of relative error
Experiment PI-DeepONet PI-DCON PI-DeepONet PI-DCON
Darcy flow (fixed mesh size) 7.10% 2.50% 4.00% 1.11%
2D plate (fixed mesh size) 5.58% 1.77% 1.14% 0.42%
Darcy flow (variable mesh size) - 3.42% - 1.43%
2D plate (variable mesh size) - 2.98% - 0.92%
In order to better visualize our model performances, we present histograms of the estimated relative error across the entire test dataset in Figures 5 and 6. Figure 5 demonstrates that, for most instances, our model predicts the PDE solutions with a relative error under 4%, whereas the relative error for most of the predictions of PI-DeepONet falls within the 4% to 10% range. In Figure 6, we also show that our model can handle varying mesh size with similar prediction accuracy as the fixed mesh size, achieving relative error under 6% for predictions of most cases. Figure 6 further illustrates that our model maintains comparable prediction accuracy across different mesh sizes. For the majority of predictions, it manages to keep the relative error below 6%.
@@
Figure 5: Comparison between PI-DCON and PI-DeepONet on the Darcy flow and 2D plate problems, with the fixed mesh size.
@@
Figure 6: Comparison of PI-DCON for fixed mesh size versus variable mesh size, on the Darcy flow and 2D plate problems.
@@
PI-DeepONet (fixed mesh size) PI-DCON (fixed mesh size) PI-DCON (variable mesh size)
@@
To further illustrate the comparative performance of our model, Tables  4 show predicted and ground truth solutions for PI-DeepONet and PI-DCON. For each problem, four of the test cases are shown. In particular, the test cases (i.e. realizations of boundary conditions) that cause the best and worst performances of PI-DeepONet and PI-DCON. It can be seen that our model shows higher accuracy, and maintains remarkable accuracy even in the most challenging (worst-case) scenario. Furthermore, the relatively small performance gap between the worst and best cases also highlights the robustness of our model. Furthermore, as can be seen in Table 5, PI-DCON can also produce acceptable predictions when training and test cases are created with variable mesh sizes.
@@
Table 5: Performance of PI-DCON trained with variable mesh sizes. The best and worst prediction among various realizations of boundary conditions are shown.

### v2

# Physics-informed Discretization-independent Deep Compositional Operator Network

@@ Abstract
Solving parametric Partial Differential Equations (PDEs) for a broad range of parameters is a critical challenge in scientific computing. To this end, neural operators, which predicts the PDE solution with variable PDE parameter inputs, have been successfully used. However, the training of neural operators typically demands large training datasets, the acquisition of which can be prohibitively expensive. To address this challenge, physics-informed training can offer a cost-effective strategy. However, current physics-informed neural operators face limitations, either in handling irregular domain shapes or in in generalizing to various discrete representations of PDE parameters. In this research, we introduce a novel physics-informed model architecture which can generalize to various discrete representations of PDE parameters and irregular domain shapes. Particularly, inspired by deep operator neural networks, our model involves a discretization-independent learning of parameter embedding repeatedly, and this parameter embedding is integrated with the response embeddings through multiple compositional layers, for more expressivity. Numerical results demonstrate the accuracy and efficiency of the proposed method.
Keywords Physics-informed neural networks, Neural operators, discretization generalization
@@ 1 Introduction
Recently, various neural operator architectures have been proposed for operator learning in a data-driven (supervised learning) approach Lu et al. (2021); Li et al. (2020a); Kovachki et al. (2021); Li et al. (2020b). As the pioneering work in this area, Deep Operator Networks (DeepONet) Lu et al. (2021) were developed, based on an inspiration from the universal approximation theorem for operators. DeepONet approximates nonlinear operators by effectively learning a collection of basis functions and coefficients, in the form of neural networks. However, the function input to DeepONet, i.e. the PDE parameters, are structured in the form of a finite-dimensional vector with a fixed dimension. This makes DeepONet to be a discretization-specific method Kovachki et al. (2021). For instance, consider employing sensors on $m$ distinct locations to observe a PDE parameter, leading to an $m$-dimensional representation of the PDE parameter. Then, a DeepONet is trained and only usable for an exactly $m$-dimensional parameter input, and won’t be applicable for cases where more or fewer sensors are used, or when discretizations with a different size are used. This restriction limits the utility of neural operators for cases where various discretizations may be used.
@@
However, in most engineering applications, we may use different mesh resolution in different location of the domain, focusing on different aspect of the system performance evaluation. To overcome the limitations of DeepONet, numerous "discretization-independent" neural operators have been introduced. These discretization-independent methods can predict the PDE solutions with various discrete representation of the PDE parameters on different meshes. The Graph Neural Operator (GNO) Li et al. (2020b) adopts graph neural networks (GNNs) architecture for operator learning, i.e. by considering graph structures for model’s input and output. The Fourier Neural Operator (FNO) Li et al. (2020a) leverages the Fourier transform to learn mappings in the spectral domain, capturing global dependencies and showing superior performance in rectangular domain shapes. Geo-FNO Li et al. (2023) extends FNO to irregular meshes by learning a mapping from an irregular mesh to a uniform mesh. Inspired by FNO, Wavelet Neural Operator (WNO) Tripura and Chakraborty (2022) uses the Wavelet Transformation, instead of the Fourier transformation, to better model signals with discontinuity and spikes, and also handle non-square domains. Also, the Low-Rank Neural Operator (LRNO) Kovachki et al. (2021) boosts the efficiency and scalability of neural operator models, using a low-rank representation of the kernels in neural operators.
@@
While these "discretization-independent" methods have shown promise in various applications, they all are data-intensive Wang et al. (2021). Training neural operators in a data-driven (supervised) way necessitates collection of a large number of parameter-solution pairs as the training set, which is obtained using repeated costly high-fidelity (FEM) simulations. This is prohibitively expensive in complex industrial applications. Under this circumstance, physics-informed neural operators emerge as an effective solution for constructing neural operators. These methods draw inspiration from physics-informed neural networks (PINNs) Raissi et al. (2019), and integrate the governing PDEs directly into the training process. This allows for a data-free training without need for FEM-based training data. Among the first physics-based methods, parametric physics-informed neural networks Zhu et al. (2023); Beltrán-Pulido et al. (2022); Naderibeni et al. (2024) considered various elements of a PDE problem, such as PDE parameters Beltrán-Pulido et al. (2022); Hansen et al. (2023), boundary conditions Zhu et al. (2023), and even the geometry of the domain Naderibeni et al. (2024), as additional feature vectors. These feature vectors provide additional inputs to the model, allowing it to learn the dependence of the PDE solution on PDE parameters. However, this approach is limited to cases where an analytical representation of varying conditions, such as boundary conditions and geometries, are available, and cannot be effective where these conditions are available in discrete form with many nodes.
@@
Another class of data-free methods are neural operators that are trained using a physics-informed loss function Goswami et al. (2023); Ren et al. (2022); Moseley et al. (2023); Baddoo et al. (2023). Among them, physics-informed DeepONet (PI-DeepONet) Wang et al. (2021) extends the original DeepONet framework by embedding physical laws directly into the loss function which can be effectively calculated in any irregular domain shape. The incorporation of physical laws also enhances the learning process but, even with this advancement, physics-informed DeepONet cannot generalize to parameter discretizations with different size. On the other hand, the physics-informed Fourier Neural Operator (PI-FNO) Li et al. (2021b) builds upon the original FNO framework, and offers generalization across different parameter dicretizations while also incorporating physics-informed training methods. However, the reliance of PI-FNO on Fast Fourier Transform (FFT) restricts its application to rectangular domains with uniform meshes as FFT works on uniform sampling of the signal in the spatial domain Li et al. (2023). Another notable development is the physics-informed Wavelet Neural Operator (PI-WNO) Navaneeth et al. (2024), which implements physics-informed training for WNO Tripura and Chakraborty (2022) using stochastic projection. Although PI-WNO can handle various PDE parameter discretizations, the hidden kernels are learned based on the architecture of convolution neural networks O’Shea and Nash (2015), which prevents the model from learning in complex domain geometries.
@@
In our study, we seek to overcome the aforementioned limitations related to data requirement, discretization-dependence and irregular geometries by introducing an innovative neural operator architecture. Our Physics-informed Deep Compositional Operator Network (PI-DCON) model is capable of generalizing across different discrete representations of PDE parameters, including those in irregular domain shapes. In this study, the irregular domain shapes are defined as those that does not belong to regular geometric forms such as rectangles, circles, or polygons. The differences between our proposed model versus other existing works are summarized in Table LABEL:table.archi_compare. To the best of our knowledge, our paper is the first attempt to develop a discretization-independent neural operator that can handle complex geometries without need for any training data.
@@ 3 Methodology
Without loss of generality, let us consider a PDE operator mapping from boundary conditions to the PDE solution, denoted by $\mathcal{M}:g(\bm{x}^{\prime})\rightarrow u(\bm{x})$. In DCON, similarly to DeepONet, the boundary conditions is encoded using a finite number of sampled boundary conditions $[\bm{x}^{\prime}_{1},\bm{x}^{\prime}_{2},...,\bm{x}^{\prime}_{m}]$, and the corresponding function values $[g(\bm{x}^{\prime}_{1}),g(\bm{x}^{\prime}_{2}),...,g(\bm{x}^{\prime}_{m})]$. The architecture of DCON consists of a "branch network" and a "solution network". The branch network maps the given boundary conditions on each of the available $m$ boundary locations to a higher dimensional space, $\mathcal{S}_{\text{param}}$, using a multi-layer perceptron. Then, a Max-pooling layer is applied to this high-dimensional representation to capture, in each dimension of $\mathcal{S}_{\text{param}}$, the most important feature across the $m$ locations on the boundary. The final output of the branch network is a global parameter embedding $\bm{b}$ in the high-dimensional space $\mathcal{S}_{\text{param}}$, whose dimension is independent of the number of boundary points, $m$. This setup relaxes the requirement for all training and test data to include the same number of boundary locations, $m$. In contrast to FNO, which aggregates spatial information after mapping spatial features to the frequency domain, we employ a simple yet effective pooling operation to directly aggregate parameter information in the original spatial domain. By avoiding the use of FFT, our model offers an efficient framework for physics-informed training using Auto Differentiation Ketkar et al. (2021).
@@
where $\sigma$ represents the nonlinear activation function, $W_{\bm{t}}\in R^{q\times d},B_{\bm{t}}\in R^{q}$ are trainable parameters of the mapping $\bm{t}$, and $W^{j}_{L}\in R^{q\times q},B^{j}_{L}\in R^{q}$ are trainable parameters of the $j$-th mapping $N$. These parameters of the model are optimized by minimizing the physics-informed training loss as shown in equation 2.
Based on the Universal Approximation Theorem, a single operator layer can approximate any continuous functional operator. Therefore, using an architecture of stacked operator layers, we can achieve at least the same approximation capability as DeepONet. Regarding model complexity, our model architecture is also comparable to DeepONet. For a hidden feature dimension of $q$, DeepONet with $L$ hidden layers has model complexity of $\mathcal{O}(Lq^{2}+q)$, considering the matrix-vector multiplication in each hidden layer and the element-wise multiplication in the last layer. In contrast, our model has similar model complexity of $\mathcal{O}(Lq^{2}+Lq)$ due to the additional element-wise multiplication in each hidden layer.
More details about the proposed model architecture are shown in Figure 2. It should be noted that the proposed DCON architecture can also be used in a data-driven approach, where its parameters are estimated in a supervised learning process, as will be discussed in Section 4.3.
@@ 4 Numerical results
In this section, we numerically evaluate the accuracy of the proposed model for solving parametric differential equations. We compare our models with data-driven DeepONet Lu et al. (2021), physics-informed DeepONet Wang et al. (2021), and an improved architecture of DeepONet reported in Wang et al. (2022). We use these default settings unless mentioned otherwise; We use the hyperbolic tangent function (Tanh) Lau and Lim (2018) as our activation function to ensure the smoothness in high-order derivatives. For the architecture of DeepONet, we use three hidden layers of width 512. For our model architecture, we use three operator layers of width 512. Comparing different learning methods is challenging because of the numerous hyper-parameters involved. To address this, we use the grid search method Bergstra and Bengio (2012) to select the appropriate hyper-parameters for each approach, ensuring the hyperparameters are properly set for different learning methods. Adam is the default optimizer with the following default hyper-parameters: $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999 Kingma and Ba (2014). Each model is trained on 70% of the sampled PDE parameters and validated on 10% of the sampled PDE parameters. Well-trained models predict the PDE solution for the remaining 20% of the sampled PDE parameters. Two hyper-parameters used in model training: (1) the learning rate, with the possible values of learning rate are 0.001, 0.0005, 0.0002, and 0.0001; and (2) the coordinate sampling size ratio in each epoch. This ratio determines the number of collocation points in each epoch as a function of the number of sampled PDE parameters. The possible values for this ratio are 0.3, 0.2, 0.1, and 0.05. Model training is performed on an NVIDIA P100 GPU using a batch size of 20. We collect the parameters of PDEs by Monte Carlo Simulation Harrison (2010) of stochastic processes Parzen (1999) and derive the solution of PDEs based on the Finite Element Method using Matlab Matlab (2012). We use the $L_{2}$ relative error ${|u_{\text{pred}}-u_{\text{fem}}|}/{|u_{\text{gt}}|}$ as our error measure, where $u_{\text{pred}}$ refers to the neural network prediction of PDE solution, and $u_{\text{fem}}$ is the PDE solutions computed by FEM.
@@ 4.1.1 A Darcy flow problem
We employ this Gaussian process model of Eq. 13 with $l=1$ to generate $N=1000$ different imposed pressure functions on $\partial\Omega_{2}$. Then given these sampled functions, $\{g_{i}\}_{i=1}^{N}$ we solve the PDE using Finite Element Method (FEM) Li and Chen (2019). In order to demonstrate the resolution-independence feature of our model, for different sampled boundary functions, we consider different mesh sizes in the domain. Specifically, for the $i$-th realization, with the boundary condition function $g_{i}$, let $M_{1,i}$ and $M_{2,i}$ be the number of sampled locations on $\partial\Omega_{1}$ and $\partial\Omega_{2}$, respectively, and $\bm{X}_{\text{BC}_{1},i}$ and $\bm{X}_{\text{BC}_{2},i}$ are the corresponding sets of coordinates at those sampled locations, i.e.
@@
To better clarify the relationship between the boundary conditions and PDE solutions, some samples of Gaussian process for boundary and PDE solution are shown in Figure 4.
Figure 4: Three representative FEM responses (bottom row) calculated for three realizations of pressure values of the boundary (top row). The dashed line represents the zero pressure.
@@ 4.1.2 A 2D plate problem
where $u$ and $v$ are the plate displacements in $x$ and $y$ directions, respectively, $E$ is the Young’s modulus, and $\mu$ is the Poisson’s Ratio.
The plate is 20 mm × 20 mm, and the hole has a diameter of 5 mm. We impose different boundary conditions on the different edges of the plate. The edge surrounding the central hole is subjected to a fixed boundary condition and there are imposed displacements on the left and right sides. The top and bottom sides are assigned free boundary conditions. The complete boundary conditions are shown in Figure  5, and are given by
@@
To better clarify the relationship between the boundary conditions and PDE solutions, some samples of Gaussian process for boundary and PDE solution are shown in Figure 6.
Figure 6: Three representative samples of the FEM responses with their corresponding prescribed displacements on the left and right edges. The dashed line represents zero displacement.
@@ 4.2 Main results
For each problem, we conduct two experiments. First, in order to compare our method with PI-DeepONet, which can only handle data with the same discrete representation, we consider a fixed discretization size, taken to be small, for all the training and test cases. In the second experiment, we relax this constraint and generate training and test data each with variable discretization sizes. The comparison between our proposed methods and the PI-DeepONet is summarized in Table  LABEL:fig.PI_com. We report the average error calculated across the entire test dataset, along with the standard deviation of the relative errors, to demonstrate the model stability. Our proposed model showed satisfactory mean performance in both the Darcy flow problem and the 2D plate problem. We found out the improved DeepONet architecture of Wang et al. (2022) to be the best baseline model. Compared to this baseline architecture, our model achieved 19.4% accuracy improvement for the Darcy flow problem and 11.9% accuracy improvement for the 2D plate problem. When the standard deviation of relative errors are considered, our model also shows higher stability.
In the second experiment, when variable variable discrete representations are used, compared to the fixed discretization case, we note that PI-DCON offers smaller accuracy levels for both the Darcy flow and the 2D plate problems. These results are reasonable, especially considering the 2D plate problem is inherently more complex. Despite these challenges, our model is able to maintain acceptable levels of accuracy, with mean relative errors below 3.5% for both problems.
Table 2: Accuracy comparison between PI-DCON and PI-DeepONet models. In the experiments with variable discrete representation, we show the model performance across different meshes.
Experiment $L_{2}$ relative error
PI-DeepONet Wang et al. (2022) PI-DCON
Darcy flow (fixed discrete representation) 7.10% $\pm$ 4.00% 3.10% $\pm$ 1.41% 2.50% $\pm$ 1.11%
2D plate (fixed discrete representation) 5.58% $\pm$ 1.14% 2.01% $\pm$ 0.53% 1.77% $\pm$ 0.42%
Darcy flow (variable discrete representation) - - 3.42% $\pm$ 1.43%
2D plate (variable discrete representation) - - 2.98% $\pm$ 0.92%
In order to better visualize our model performances, we present histograms of the estimated relative error across the entire test dataset in Figures 7 and 8. Figure 7 demonstrates that, for most instances, our model predicts the PDE solutions with a relative error under 4%, whereas the relative error for most of the predictions of PI-DeepONet falls within the 4% to 10% range. Compared with Wang et al. (2022), our model also show slightly higher performance. In Figure 8, we also show that our model can handle different discretization sizes with similar prediction accuracy as the fixed discretization size, achieving relative error under 6% for predictions of most cases. Figure 8 further illustrates that our model maintains comparable prediction accuracy across different discretization sizes. For the majority of predictions, it manages to keep the relative error below 6%.
@@
Figure 7: Comparison between PI-DCON and PI-DeepONet on the Darcy flow and 2D plate problems, with the fixed discretization size.
@@
Figure 8: Comparison of PI-DCON for fixed discretization size versus variable discretization size, on the Darcy flow and 2D plate problems.
@@
PI-DeepONet (fixed discrete representation) PI-DCON (fixed discrete representation) PI-DCON (variable discrete representation)
@@
To further illustrate the comparative performance of our model, Tables  4 show predicted and ground truth solutions for PI-DeepONet and PI-DCON. For each problem, four of the test cases are shown. In particular, the test cases (i.e. realizations of boundary conditions) that cause the best and worst performances of PI-DeepONet and PI-DCON. It can be seen that our model shows higher accuracy, and maintains remarkable accuracy even in the most challenging (worst-case) scenario. Furthermore, the relatively small performance gap between the worst and best cases also highlights the robustness of our model. Furthermore, as can be seen in Table 5, PI-DCON can also produce acceptable predictions when training and test cases are created with variable discretization sizes.
@@
Table 5: Performance of PI-DCON trained with variable discrete representations. The best and worst prediction among various realizations of boundary conditions are shown.